OpenStudy (anonymous):

need help with the set up...find two consecutive odd integers such that the sum of the smaller and 4 times the larger is 33.

OpenStudy (anonymous):

x+4(x+2)>=33,

OpenStudy (anonymous):

x - y = 2 4x + y = 33 add 5x = 35 x = 7 so y = 5 answer 5,7

OpenStudy (anonymous):

then I distribute the 4 first?

OpenStudy (anonymous):

x+4(x+2)>=33, 5x+8>=33, 5x=25, x=5, y=7

OpenStudy (anonymous):

great thank you!